OpenStudy (narissa):

Please help have to spend the rest of my weekend catching up... student solved 6x+4=2(x−8) in four steps as shown: Step 1: 6x+4=2x−8 Step 2: 4x+4=−8 Step 3: 4x=−12 Step 4: x=−3 Which statement is an accurate interpretation of the students work? Question 1 options: The correct solution is x=−3 The student made an error in Step 1. The student made an error in Step 3.

OpenStudy (narissa):

@roadjester

OpenStudy (a1234):

What do you think @narissa ?

OpenStudy (a1234):

No. Look carefully if the student correctly used the distributive property.

OpenStudy (narissa):

oh B

OpenStudy (a1234):

Can you tell me how you got that answer?

OpenStudy (narissa):

i guessed... im kinda slow in math

OpenStudy (narissa):

sorry my computer froze

OpenStudy (a1234):

Look at this website and compare it to the steps you are given. http://www.coolmath.com/prealgebra/06-properties/05-properties-distributive-01

OpenStudy (narissa):

he didnt multiply the numbers correctly

OpenStudy (roadjester):

I agree with @a1234 The distributive law was incorrectly applied.

OpenStudy (narissa):

so b is the correct answer..?

OpenStudy (a1234):

Yes, b is the correct answer.

OpenStudy (anonymous):

Add 8 on both sides,so that the equation will be 6x+12 = 2x. Then divide both sides by 2 so that the equation will be 3x+6 = x. Divide x on both sides so that you get 3+6 = x. S0 x = 9. That or do distributive property on the right side so that this is how the equation looks like: 6x+4 = 2x-16. Then add 16 on both sides, now you have 6x+4+16 = 2x. Rewrite the equation as 6x+20 = 2x.Then divide both sides by 2 so that you get: 3x+10 = x divide by x on both sides so that its 3+10 = x. So x = 13, so depending on how you do the problem the x could equal either 9 or 13.

OpenStudy (a1234):

@StudiousHojea I see you want to help, but dividing both sides by x does not work (and it's not relevant to the problem).

OpenStudy (anonymous):

How does it not work and how is it not relevant? I am not annoyed or anything, just curious.

OpenStudy (a1234):

Because you are trying to find what the variable is. If you divide out the variable, you're not really getting any results. When you solve an equation you try to separate the variable from the constants. So... 6x + 4 = 2x - 16 4x = -20 x = -5

OpenStudy (a1234):

Actually, if you divide out the variable in 6x + 4 = 2x - 16, it does work, but it has to be done correctly. That would be 6 + 4x = 2 - 16x 4 = -20x x = -5 So my response above was partly incorrect

OpenStudy (anonymous):

Oh! Now get it. So after writing 6x+4 = 2x-16, you have to subtract 4 from both sides so that you get 6x = 2x-20. Then you subtract 2x from both sides so its 4x = -20 and that is how you get x = -5?

OpenStudy (a1234):

Correct!

OpenStudy (anonymous):

Wow! Thank you! I think I understand math better a little.

OpenStudy (a1234):

You're welcome. And good questions!

OpenStudy (anonymous):

Thanks! Have a nice day! Well, I think we are going to talk again sometime today, but for now, have a good day!

OpenStudy (anonymous):

What grade are you in? I am in 10th